### v1

# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

Abstract
Signal Temporal Logic (STL) is a powerful specification language for describing complex temporal behaviors of continuous signals, making it well-suited for high-level robotic task descriptions. However, generating executable plans for STL tasks is challenging, as it requires consideration of the coupling between the task specification and the system dynamics. Existing approaches either follow a model-based setting that explicitly requires knowledge of the system dynamics or adopt a task-oriented data-driven approach to learn plans for specific tasks. In this work, we investigate the problem of generating executable STL plans for systems whose dynamics are unknown a priori. We propose a new planning framework that uses only task-agnostic data during the offline training stage, enabling zero-shot generalization to new STL tasks. Our framework is hierarchical, involving: (i) decomposing the STL task into a set of progress and time constraints, (ii) searching for time-aware waypoints guided by task-agnostic data, and (iii) generating trajectories using a pre-trained safe diffusion model. Simulation results demonstrate the effectiveness of our method indeed in achieving zero-shot generalization to various STL tasks.
1 Introduction
Signal Temporal Logic (STL) is a formal specification language used to describe the temporal behavior of continuous signals. It has become widely adopted for specifying high-level robotic behaviors due to its expressiveness and the availability of both Boolean and quantitative evaluation measures. Controlling robots under STL task constraints, however, is a challenging problem, as it requires balancing both the satisfaction of the task and the feasibility of the system dynamics. In cases where the environment and system dynamics are fully known, several representative methods have been developed, including optimization-based approaches Raman et al. (2014); Kurtz and Lin (2022); Sun et al. (2022), gradient-based techniques Gilpin et al. (2020); Dawson and Fan (2022), and sampling-based methods Ilyes et al. (2023). However, these methods are often difficult to apply in practical scenarios, where the system dynamics and environment are either unknown or difficult to model.
To address the challenge of unknown dynamics, several learning-based approaches have been proposed. One typical method is reinforcement learning (RL) Aksaray et al. (2016); Balakrishnan and Deshmukh (2019); Kalagarla et al. (2021); Venkataraman et al. (2020); Ikemoto and Ushio (2022); Wang et al. (2024), where an appropriate reward function is designed to approximate the satisfaction of the STL task. However, these methods often struggle with long-horizon STL tasks and lack generalization capabilities across different tasks. Another approach involves first learning a system model and then integrating it with model-based planning methods. For example, in Kapoor et al. (2020), the authors trained a neural network to approximate the system dynamics and combined it with an optimization-based approach. However, this method is limited to simple short-horizon STL tasks due to its high computational cost. In He et al. (2024), the authors used goal-conditioned RL to train multiple goal-conditioned policies, referred to as “skills,” to accomplish specific objectives. They then applied a search algorithm to determine the optimal sequence of “skills” needed to satisfy the given STL tasks. While this approach enables a certain degree of task generalizations, these tasks must be based on pre-defined objectives associated with the skills.
More recently, generative models, such as diffusion models Ho et al. (2020), have emerged as a new approach for generating trajectories for systems with unknown dynamics Janner et al. (2022); Ajay et al. (2022); Chi et al. (2023); Carvalho et al. (2023); Huang et al. (2025), gaining popularity across various applications. Compared to traditional model-based reinforcement learning methods, these generative approaches are better suited for long-horizon decision-making and offer greater test-time flexibility Janner et al. (2022), making them particularly effective for complex tasks. For example, for finite Linear Temporal Logic ($\text{LTL}_{f}$) tasks, Feng et al. (2024a) introduced a classifier-based guidance approach to steer the sampling of diffusion models, ensuring that generated trajectories satisfy $\text{LTL}_{f}$ requirements. Similarly, Feng et al. (2024b) proposed a hierarchical framework that decomposes co-safe LTL tasks into sub-tasks using hierarchical reinforcement learning. This framework employs a diffusion model with a determinant-based sampling strategy to generate diverse low-level trajectories, improving both planning success rates and task generalization.
In the context of STL trajectory planning, the use of generative models has also been explored recently. For example, Zhong et al. (2023) proposed a classifier-based guidance approach that leverages robustness gradients to guide diffusion model sampling, enabling the generation of vehicle trajectories that adhere to traffic rules specified by STL. Building on this, Meng and Fan (2024) introduced a data augmentation method to enhance trajectory diversity and improve rule satisfaction rates. However, these approaches are still limited to simpler STL tasks, primarily due to the complexity of optimizing robustness values and the inherent trade-off between maximizing reward objectives and maintaining the feasibility of the generated trajectories Li et al. (2024).
In this paper, we address the challenge of generating trajectories for complex, long-horizon STL tasks that are feasible for an underlying system with unknown dynamics. Specifically, we assume access only to a set of task-agnostic trajectory data from previous operations. Inspired by recent advances in decomposition-based STL planningKapoor et al. (2024), we propose a novel hierarchical framework that integrates task decomposition, search algorithms, and generative models. First, complex STL tasks are decomposed into a set of time-aware reach-avoid progresses and time constraints. Next, a search algorithm, heuristically guided by the trajectory data, is employed to allocate these progresses and generate a sequence of waypoints with corresponding timestamps. Finally, a pre-trained diffusion model, trained on task-agnostic data, is used to sequentially generate trajectories that achieve the timed waypoints, resulting in a complete solution. To the best of our knowledge, our algorithm is the first data-driven approach with zero-shot generalization capabilities for complex STL tasks. Simulation experiments demonstrate that our method achieves a high success rate in trajectory planning across diverse, long-horizon STL tasks and outperforms commonly used non-data-driven methods in terms of efficiency.
2 Preliminaries
2.1 System Model
We consider a discrete time system with unknown dynamics
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{a}_{t}),$ (1)
where $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and $\mathbf{a}_{t}\in\mathbb{R}^{m}$ are the state and the action at time instant $t$, respectively. Given an initial state $\mathbf{x}_{0}$ and a sequence of actions $\mathbf{a}_{0}\mathbf{a}_{1}\dots\mathbf{a}_{T-1}$, the resulting trajectory of the system is $\boldsymbol{\tau}=\mathbf{x}_{0}\mathbf{a}_{0}\mathbf{x}_{1}\mathbf{a}_{1}\dots\mathbf{a}_{T-1}\mathbf{x}_{T}$, where $T$ is the horizon. The signal of the trajectory is referred to as the state sequence $\mathbf{s}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$ and we denote by $\mathbf{s}_{t}=\mathbf{x}_{t}\mathbf{x}_{t+1}\dots\mathbf{x}_{T}$ the sub-signal starting from time step $t$.
2.2 Signal Temporal Logic
We use signal temporal logic (STL) to describe the formal task imposed on the generated state sequence Maler and Nickovic (2004). Specifically, we consider STL formula in the Positive Normal Form (PNF) Sadraddini and Belta (2015) whose syntax is as follows:
$\varphi\!::=\!\top\mid\mu\mid\varphi_{1}\!\wedge\!\varphi_{2}\mid\varphi_{1}\!\vee\!\varphi_{2}\mid\mathrm{F}_{[a,b]}\varphi\mid\mathrm{G}_{[a,b]}\varphi\mid\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2},$ (2)
where $\top$ is the true predicate and $\mu$ is an atomic predicate associated with an evaluation function $h_{\mu}:\mathbb{R}^{n}\to\mathbb{R}$, i.e., predicate $\mu$ is true at state $\mathbf{x}_{t}$ iff $h_{\mu}(\mathbf{x}_{t})\geq 0$. Furthermore, $\wedge$ and $\vee$ are logic operators “conjunction” and “disjunction”, respectively; $\mathrm{U}_{[a,b]},\mathrm{F}_{[a,b]}$ and $\mathrm{G}_{[a,b]}$ are temporal operators “until”, “eventually” and “always”, respectively; $[a,b]$ is a time interval such that $a,b\in\mathbb{Z},0\leq a\leq b<\infty$. Note that, negation is not used in the PNF. However, as shown in Sadraddini and Belta (2015), this does not result in any loss of generality as one can always redefine atomic predicates to account for the presence of negations, allowing any general STL formula to be expressed in PNF. In our work, we impose an additional restriction on the Prenex Normal Form of formulas. Specifically, for any formula of the form $\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2}$, $\varphi_{1}$ can only involve temporal operator “always”. This restriction is introduced for technical reasons, as it facilitates the decomposition of the overall formula into a set of progresses.
For any signal $\mathbf{s}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$, we denote by $\mathbf{s}_{t}\vDash\varphi$ if $\mathbf{s}$ satisfies STL formula $\varphi$ at time $t$, and we denote by $\mathbf{s}\vDash\varphi$ if $\mathbf{s}_{0}\vDash\varphi$. This is formal defined by the Boolean semantics of STL formulae as follows Bartocci et al. (2018):
$\mathbf{s}_{t}\vDash\mu$ $\Leftrightarrow h_{\mu}(\mathbf{x}_{t})\geq 0,$ (3)
$\mathbf{s}_{t}\vDash\varphi_{1}\wedge\varphi_{2}$ $\Leftrightarrow\mathbf{s}_{t}\vDash\varphi_{1}\wedge\mathbf{s}_{t}\vDash\varphi_{2},$ (4)
$\mathbf{s}_{t}\vDash\varphi_{1}\vee\varphi_{2}$ $\Leftrightarrow\mathbf{s}_{t}\vDash\varphi_{1}\vee\mathbf{s}_{t}\vDash\varphi_{2},$ (5)
$\mathbf{s}_{t}\vDash\mathrm{F}_{[a,b]}\varphi$ $\Leftrightarrow\exists t^{\prime}\in[t+a,t+b]\text{ s.t. }\mathbf{s}_{t^{\prime}}\vDash\varphi,$ (6)
$\mathbf{s}_{t}\vDash\mathrm{G}_{[a,b]}\varphi$ $\Leftrightarrow\forall t^{\prime}\in[t+a,t+b]\text{ s.t. }\mathbf{s}_{t^{\prime}}\vDash\varphi,$ (7)
$\mathbf{s}_{t}\vDash\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2}$ $\Leftrightarrow\exists t^{\prime}\in[t+a,t+b]\text{ s.t. }\mathbf{s}_{t^{\prime}}\vDash\varphi_{2}$
$\quad\wedge\forall t^{\prime\prime}\in\left[t,t^{\prime}\right],\mathbf{s}_{t^{\prime\prime}}\vDash\varphi_{1}.$ (8)
2.3 Planning with Unknown Dynamics
In the context of STL planning, the objective is to determine an action sequence such that the resulting signal satisfies the specified STL formula. When the system dynamics are perfectly known, this problem can be solved using model-based optimization approaches (e.g., see Raman et al. (2014); Kurtz and Lin (2022); Sun et al. (2022)). In contrast, our work addresses a setting with unknown dynamics. Specifically, we assume the mapping $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is unknown,but a dataset of historical operational trajectories, consistent with the underlying unknown system dynamics, is available. Note that each trajectory in the dataset is collected from the previous task-agnostic operations and may vary in length. Our goal is to leverage these task-agnostic trajectories to generate new trajectories that satisfy any given STL formula, thereby achieving zero-shot task generalization at test time.
Problem 1.
Given a set of trajectories from the unknown system (1) and a STL formula $\varphi$, find a sequence of actions $\mathbf{a}_{0}\mathbf{a}_{1}\dots\mathbf{a}_{T}$ such that the resulting signal $\mathbf{s}$ satisfies the STL formula, i.e., $\mathbf{s}\vDash\varphi$.
3 Our Method
3.1 Overall Framework
Figure 1: The Overall Framework of Our Proposed Method
First, we provide an overview of our proposed planning methods, whose overall structure is illustrated in Figure 1. Specifically, our method consists of the following three parts:
  • Task Decomposition: First, we decompose the given STL into a set of spatial-temporal relevant progresses $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}^{\mathcal{I}}$, where $\mathbb{P}^{\mathcal{R}}$ is the set of reachability progresses and $\mathbb{P}^{\mathcal{I}}$ is the set of invariance progresses. Furthermore, these progresses need to be satisfied subject to a set of time constraints $\mathbb{T}$ over a set of time variables $\Lambda$.
  • Progress Allocation: While the decomposition above is task-focused, achieving progress in the correct order requires consideration of the underlying dynamics of the system. To address this, we employ a pre-trained time predictor derived from trajectory data to estimate the time required for the system to transition from one waypoint to another. A search-based algorithm is then used to determine a sequence of waypoints with associated timestamps $(\mathbf{x}_{0},t_{0})(\mathbf{x}_{1},t_{1})\dots(\mathbf{x}_{n},t_{n})$ such that each waypoint satisfies the corresponding reachability progress in $\mathbb{P}^{\mathcal{R}}$, and all timestamps comply with the time constraints $\mathbb{T}$.
  • Trajectory Generation: The final step is to generate an executable trajectory for the system, ensuring that each waypoint is visited at the correct time while maintaining satisfaction of the invariance progresses. This problem can be framed as several reach-avoid control problems with unknown system dynamics. Our approach leverages a diffusion model, pre-trained on trajectory data, to generate task-compliant and dynamic-feasible trajectories through constraints-guided sampling. Following Ajay et al. (2022), we utilize the diffusion model solely for generating the trajectory’s state sequence. The action sequence can be obtained offline using an inverse dynamics model Agrawal et al. (2016) or online via appropriate controllers.
Nexet, we provide the technical details of each part.
3.2 Decompositions of STL Formulae
Eliminating Disjunctions
For the given STL task $\varphi$, our first step is to covert it into the disjunctive normal form (DNF) $\tilde{\varphi}=\varphi_{1}\vee\varphi_{2}\vee...\vee\varphi_{n}$, where each subformula $\varphi_{i}$ involves no “disjunction”. Formally, for any STL formula, one can obtain its DNF by recursively applying the following replacements:
  • replace $\mathrm{F}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ by $\mathrm{F}_{[a,b]}\varphi_{1}\vee\mathrm{F}_{[a,b]}\varphi_{2}$;
  • replace $\mathrm{G}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ by $\mathrm{G}_{[a,b]}\varphi_{1}\vee\mathrm{G}_{[a,b]}\varphi_{2}$;
  • replace $(\phi_{1}\!\vee\!\phi_{2})\mathrm{U}_{[a,b]}(\varphi_{1}\!\vee\!\varphi_{2})$ by $\bigvee_{i,j\in\{1,2\}}\phi_{i}\mathrm{U}_{[a,b]}\varphi_{j}$.
Note that the last two replacements are not equivalent, making the resulting DNF $\tilde{\varphi}$ stronger than the original formula ${\varphi}$. Furthermore, to achieve the task defined by $\tilde{\varphi}$, it suffices to satisfy one of the subformulas $\varphi_{i}$. Without loss of generality, we will assume henceforth that the DNF contains only a single subformula, as the STL planning problem can be addressed for each subformula individually. In other words, moving forward, we will focus on STL formulae, denoted directly by $\varphi$, without negations (due to the PNF) and without disjunctions (due to the DNF).
Progresses and Constraints
Next, we further decompose the overall STL task $\varphi$ into a set of progresses. Specifically, we consider the following two types of progresses:
  • Reachability Progress: we denote by $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ that $\exists t\in[a_{\Lambda},b_{\Lambda}],\mathbf{x}_{t}\vDash\mu$.
  • Invariance Progress: we denoted by $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$ that $\forall t\in[a_{\Lambda},b_{\Lambda}],\mathbf{x}_{t}\vDash\mu$.
Note that we use subscript ${\Lambda}$ in time interval $[a_{\Lambda},b_{\Lambda}]$ as $a_{\Lambda}$ and $b_{\Lambda}$ may not be a fixed time and involve variables in $\Lambda$ subject to time constraints. Therefore, the STL formula $\varphi$ is decomposed into a tuple $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$, where $\mathbb{P}_{\varphi}$ is the set of progresses and $\mathbb{T}_{\varphi}$ is the set of time constraints over variable set $\Lambda$. Such decomposition is defined recursively as follows:
  • If $\varphi=\mathrm{F}_{[a,b]}\mu$, then we have $\mathbb{P}_{\varphi}=\{\mathcal{R}(\lambda_{i},\lambda_{i},\mu)\}$ and $\mathbb{T}_{\varphi}=\{\lambda_{i}\in[a,b]\}$, where $\lambda_{i}$ is a new time variable.
  • If $\varphi=\mathrm{G}_{[a,b]}\mu$, then we have $\mathbb{P}_{\varphi}\!=\!\{\mathcal{I}(a,b,\mu)\}$ and $\mathbb{T}_{\varphi}\!=\!\emptyset$.
  • If $\varphi=\mu_{1}\mathrm{U}_{[a,b]}\mu_{2}$, then we have $\mathbb{P}_{\varphi}\!=\!\{\mathcal{I}(a,\lambda_{i},\mu_{1}),$ $\mathcal{R}(\lambda_{i},\lambda_{i},\mu_{2})\}$ and $\mathbb{T}_{\varphi}=\{\lambda_{i}\in[a,b]\}$, where $\lambda_{i}$ is a new time variable.
  • If $\varphi^{\prime}=\varphi_{1}\wedge\varphi_{2}$, then we merge the progresses and time constraints, i.e., $\mathbb{P}_{\varphi}=\mathbb{P}_{\varphi_{1}}\cup\mathbb{P}_{\varphi_{2}}$ and $\mathbb{T}_{\varphi}=\mathbb{T}_{\varphi_{1}}\cup\mathbb{T}_{\varphi_{2}}$.
  • If $\varphi^{\prime}=\mathrm{F}_{[a,b]}\varphi$, then we (i) introduce a new time variable $\lambda_{i}$; (ii) add a new time constraint $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}\cup\{\lambda_{i}\!\in\![a,b]\}$; (iii) increase each time indices in each progress by $\lambda_{i}$, i.e., $\mathbb{P}_{\varphi^{\prime}}=\{\mathcal{P}(c_{\Lambda}+\lambda_{i},d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{P}(c_{\Lambda},d_{\Lambda},\mu)\!\in\!\mathbb{P}_{\varphi}\}$.
  • If $\varphi^{\prime}=\mathrm{G}_{[a,b]}\varphi$, then (i) the time constraints remain unchanged, i.e., $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}$; (ii) modify each invariance progress $\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$ to $\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+b,\mu)$ in $\mathbb{P}_{\varphi^{\prime}}$; (iii) modify each reachability progress $\mathcal{R}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$, to $b-a$ progresses $\mathcal{R}(c_{\Lambda}+k,d_{\Lambda}+k,\mu)$ in $\mathbb{P}_{\varphi^{\prime}}$, where $k=a,a+1,\dots,b$.
  • If $\varphi^{\prime}=\phi\mathrm{U}_{[a,b]}\varphi$, then (i) introduce a new time variable $\lambda_{i}$; and (ii) add a new time constraint, i.e., $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}\cup\{\lambda_{i}\in[a,b]\}$; and (iii) increase each time indices in each progress in $\mathbb{P}_{\varphi}$ by $\lambda_{i}$ and modify each invariance progress $\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\phi}$ to $\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+\lambda_{i},\mu)$, i.e., $\mathbb{P}_{\varphi^{\prime}}=\{\mathcal{P}(c_{\Lambda}+\lambda_{i},d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{P}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}\}\cup\{\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\phi}\}$.
To illustrate the above progress decomposition process, we consider the following STL formulae
$\varphi=\mathrm{F}_{[5,12]}(\mathrm{F}_{[7,16]}\mu_{1}\wedge\mathrm{G}_{[2,10]}\mu_{2})\wedge\mathrm{G}_{[18,20]}\mathrm{F}_{[4,10]}\mu_{3}.$ (9)
The decomposition process is shown in Figure 2, where the progress and time constraints are constructed incrementally from the bottom to the top. The top node represents the overall decomposed $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$ for the STL formula $\varphi$.
Figure 2: Decomposition Process of STL Formula (9)
3.3 Progress Allocation
Without considering the system dynamics, the STL task planning problem is essentially a constraint satisfaction problem for the decomposed progresses and constraints $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$. However, the unknown system dynamics introduce additional challenges, as allocating progress arbitrarily may not be feasible for the system. To address this issue, we use a search algorithm to allocate the progresses, where the feasibility of each assignment is determined based on the trajectory data.
To perform the search-based allocation process, we further split the progresses $\mathbb{P}_{\varphi}$ as follows. For each invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$, we decompose it into a reachability progress $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and an invariance progress $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. For simplicity, we will denote the further decomposed progresses as $(\mathbb{P},\mathbb{T})$ without subscripts, where $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}^{\mathcal{I}}$. Due to this further decomposition, each invariance progress follows a unique reachability progress.
Main Allocation Algorithm
Algorithm 1 Main-Allocation
0:  Initial state $\mathbf{x}_{0}$, start time $t_{0}$, reachability progresses $\mathbb{P}^{\mathcal{R}}$, invariance progresses $\mathbb{P}^{\mathcal{I}}$, time variable constraints $\mathbb{T}$
0:  A valid waypoints sequence $\bar{\mathbf{s}}$ or None if no solution is found
1:  Initialize:
2:  current state $\mathbf{x}\leftarrow\mathbf{x}_{0}$; current time $t\leftarrow t_{0}$
3:  task sequence $\bar{\mathbf{s}}\leftarrow[(\mathbf{x},t)]$
4:  stack $\leftarrow[(\mathbf{x},t,\mathbb{P}^{\mathcal{R}},\mathbb{T},\bar{\mathbf{s}})]$
5:  while stack is not empty do
6:   $(\mathbf{x},t,\mathbb{P}^{\mathcal{R}},\mathbb{T},\bar{\mathbf{s}})\leftarrow$ pop(stack)
7:   if $\mathbb{P}^{\mathcal{R}}=\emptyset$ then
8:    return $\bar{\mathbf{s}},\mathbb{T}$  // All reachability progresses satisfied
9:   for each progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{R}}$ do
10:    $t^{\prime},\mathbf{x}^{\prime}\leftarrow\texttt{SampleState}(\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu),\mathbf{x},t,\mathbb{T},\mathbb{P}^{\mathcal{I}})$
11:    if $t^{\prime}\neq\texttt{None}$ then
12:     $\bar{\mathbf{s}}^{\prime}\leftarrow\bar{\mathbf{s}}.(\mathbf{x}^{\prime},t^{\prime})$
13:     ${\mathbb{P}^{\mathcal{R}}}^{\prime}\leftarrow\mathbb{P}^{\mathcal{R}}\setminus\{\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)\}$
14:     ${\mathbb{T}}^{\prime}\leftarrow\texttt{UpdateConstraint}(a_{\Lambda},b_{\Lambda},\mathbb{T},\mathbf{x}^{\prime},t^{\prime})$
15:     // Push new state onto the stack
16:     push $\bigl{(}\mathbf{x}^{\prime},t^{\prime},{\mathbb{P}^{\mathcal{R}}}^{\prime},{\mathbb{T}}^{\prime},\bar{\mathbf{s}}^{\prime}\bigr{)}$ onto stack
17:   end for
18:  end while
19:  return None  // No valid sequence found
The main algorithm for progress allocation is presented in Algorithm 1. The algorithm employs a depth-first search (DFS) to sequentially assign satisfaction times and waypoint for each reachability progress in $\mathbb{P}^{\mathcal{R}}$. When the algorithm terminates, it returns a sequence of waypoints with associated timestamps of form $(x_{0},t_{0})(x_{1},t_{1})\dots(x_{n},t_{n})$ such that each waypoint corresponds to the satisfication of a reachability progress in $\mathbb{P}^{\mathcal{R}}$. During the search process, we maintain the current state $\mathbf{x}$, the current time step $t$, the set of remaining reachability progresses $\mathbb{P}^{\mathcal{R}}$, the set of all time constraints $\mathbb{T}$, and the searched sequence of waypoints with associated timestamps $\bar{\mathbf{s}}$.
At each step, we select a remained reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$ as the next target progress from the current state $(\mathbf{x},t)$. Specifically, to achieve progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$, function SampleState is used to determine a satisfaction time $t^{\prime}$ and a corresponding new state $\mathbf{x}^{\prime}$ such that $\mathbf{x}^{\prime}\models\mu$. Then the searched waypoint $(\mathbf{x}^{\prime},t^{\prime})$ is appended to $\bar{\mathbf{s}}$ and we remove progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$. Furthermore, the time constraints $\mathbb{T}$ are updated based on the assigned state $\mathbf{x}^{\prime}$ and time $t^{\prime}$ according to function UpdateConstraint. Finally, the algorithm proceeds to the next iteration. If $\mathbb{P}^{\mathcal{R}}$ becomes empty, it indicates that all reachability progresses have been successfully assigned a satisfaction time. If no feasible assignment can be made, the algorithm backtracks to explore alternative assignments.
In the above DFS, we use the following heuristic order to select reachability progress from $\mathbb{P}^{\mathcal{R}}$. Let $\mathbb{T}$ be a set of time constraints and $a_{\Lambda}$ be an objective function over time variables $\Lambda$. We denote by $a_{\Lambda,\mathbb{T}}^{\mathrm{max}}$ and $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ the maximum and minimum values of $a_{\Lambda}$ under $\mathbb{T}$, respectively, which can be easily solved by Integer Linear Programming (ILP) techniques. Then during the search process, progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ with earlier potential deadlines $b_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ are prioritized; and if two progresses have the same deadline, the one with the earlier potential start time $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ is preferred.
Constraint Update
Let $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ be the selected progress and $(\mathbf{x}^{\prime},t^{\prime})$ be the assigned waypoint with time. Then we added the following time constraints to $\mathbb{T}$:
$a_{\Lambda}\leq t^{\prime}\quad\text{ and }\quad t^{\prime}\leq b_{\Lambda}.$ (10)
Recall that in our decomposition, each original invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$ is decomposed to $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. Therefore, if constraints (10) are added for $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$, it means that the value of $a_{\Lambda}$ is determined as $t^{\prime}$, and we denoted by $\mathbb{P}^{\mathcal{I}}_{det}$ the set of invariance progresses whose starting times are determined. Then for those invariance progress $\mathcal{I}(c,d_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}}_{det}$ whose starting time is determined, if $x^{\prime}\nvDash\mu$, then we further add a new constraint $d_{\Lambda}<t^{\prime}$, i.e., the invariance progress can only be effective before $t^{\prime}$ to avoid conflict with the assigned waypoint.
Sample Timed Waypoints
Algorithm 2 SampleState
0:  reachability progress $\mathcal{R}(a,b,\mu)$, current state $\mathbf{x}$, time step $t$, time variable constraints $\mathbb{T}$, invariance progresses $\mathbb{P}^{\mathcal{I}}$
0:  Assigned satisfaction time $t_{\mathrm{new}}$ of constraint $\mathcal{R}(a,b,\mu)$ and new state $\mathbf{x}^{\prime}$ or None if no solution is found
1:  $t_{\mathrm{min}}\leftarrow a_{\Lambda,\mathbb{T}}^{\mathrm{min}},t_{\mathrm{max}}\leftarrow b_{\Lambda,\mathbb{T}}^{\mathrm{max}}$
2:  for up to $N_{\mathrm{max}}$ attempts do
3:     Sample state $\mathbf{x}^{\prime}$ such that $\mathbf{x}^{\prime}\models\mu$
4:     Initialize: conflict time interval $\mathcal{O}\leftarrow\emptyset$
5:     for all $\mathcal{I}(c,d_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}}_{det}$ with determined starting time do
6:        if $\mathbf{x}^{\prime}\nvDash\mu$ then
7:           $\mathcal{O}\leftarrow\mathcal{O}\cup[c,d_{\Lambda,\mathbb{T}}^{\mathrm{min}}]$
8:        end if
9:     end for
10:     $t^{\prime}\leftarrow t+\texttt{TimePredict}(\mathbf{x},\mathbf{x}^{\prime})$
11:     if $t^{\prime}>t_{\mathrm{max}}$ or $[\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}=\emptyset$ then
12:        Continue to next sampling attempt
13:     end if
14:     $t_{\mathrm{new}}\leftarrow$earliest time in $[\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}$
15:     return  $t_{\mathrm{new}},\mathbf{x}^{\prime}$
16:  end for
17:  return  None  // No valid time found
When a reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ is selected, we use function SampleState to determine a valid satisfaction time and waypoint state for this progress while ensuring compliance with invariance progresses. The pseudocode of this fucntion is shown in Algorithm 2. The process starts by computing the largest possible time interval $[t_{\mathrm{min}},t_{\mathrm{max}}]$ for the reachability progress. Then the algorithm attempts to sample a candidate state $\mathbf{x}^{\prime}$ with the satisfication region of $\mu$ up to $N_{\mathrm{max}}$ times. For each attempt, we first calculate the minimum possible conflict time interval, denoted by $\mathcal{O}$, during which $\mathbf{x}^{\prime}$ conflicts with the invariance progresses. Here, we only consider invariance progresses whose starting times are determined. This is because, in the STL decomposition, each invariance progress follows a “preceding” reachability progress. If the starting time of an invariance progress is not determined, then it implies that its “preceding” reachability progress has not yet been satisfied and will only be satisfied strictly later than the current reachability progress. Consequently, this invariance progress will also start strictly later.
Once the conflict time interval $\mathcal{O}$ is computed, we further use function TimePredict to predict the arrival time $t^{\prime}$ from the current state $\mathbf{x}$ to the sample state $\mathbf{x}^{\prime}$. Particularly, if (i) $t^{\prime}>t_{\mathrm{max}}$; or (ii) the feasible interval $[\max(t^{\prime},t_{\mathrm{min}}),t_{\mathrm{max}}]$ is fully occupied by conflicting intervals in $\mathcal{O}$, then it means that the sample state $\mathbf{x}^{\prime}$ is not feasible and we proceed to the next attempt. Otherwise, the earliest available time $t_{\mathrm{new}}$ in the feasible interval is assigned as the satisfaction time, and the algorithm returns $t_{\mathrm{new}}$ along with the sampled state $\mathbf{x}^{\prime}$.
Prediction of Reachability Time
In Algorithm 2, model TimePredict is used to estimates the time step (trajectory length) needed to transition from the current state $\mathbf{x}$ to the new state $\mathbf{x}^{\prime}$. This model is trained on the same trajectory data, that will be used to train the Diffusion model. It assumes that the trajectory length between two states follows a Gaussian distribution. A simple multilayer perceptron (MLP) is used as the backbone of TimePredict model, and it is trained to predict the mean and variance of the trajectory length using a negative log-likelihood loss function. To account for tasks with avoidance requirements, which may require longer trajectories, a scaling factor $\gamma$ is applied to the predicted mean trajectory length. This factor allows control over the algorithm’s conservativeness by adjusting the predicted trajectory length as needed.
3.4 Trajectory Generation
In the above part, a sequence of waypoints $\bar{\mathbf{s}}$ is obtained and the time intervals for each invariance progress have also been determined. The remaining task is to generate executable trajectories that connect the waypoints while ensuring the trajectories satisfy the invariance progresses. Specifically, the trajectory $\tau$ between two adjacent waypoints $(\mathbf{x}_{i},t_{i})$ and $(\mathbf{x}_{i+1},t_{i+1})$ must be feasible under the current system dynamics and satisfy the following conditions:
  1. 1.
    The trajectory length is $t_{i+1}-t_{i}+1$.
  2. 2.
    It starts at state $\mathbf{x}_{i}$ and ends at state $\mathbf{x}_{i+1}$.
  3. 3.
    It satisfies all invariance progresses active within the time interval $[t_{i},t_{i+1}]$.
We employ the diffusion model to solve this conditional trajectory generation problem since it can learn the distribution of system trajectories from system trajectory data, enabling it to generate feasible trajectories under the current system. Additionally, by guiding the sampling process, the generated trajectories can satisfy additional constraints.
The technical details of the diffusion model employed are provided in the Appendix B. Here, we briefly outline the main idea. To satisfy the first condition, we control the length of the generated trajectory by adjusting the length of the initial noise during the denoising process. This is feasible due to the structural properties of the backbone network used in the diffusion model Janner et al. (2022). For the second condition, we treat it as an inpainting problem Janner et al. (2022), ensuring that the generated trajectory meets the requirement by replacing the start and end states of the trajectory with $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$, respectively, after each denoising step. To address the last condition, we decompose each invariance progress $\mathcal{I}(a,b,\mu_{i})$ into state constraints of the form $h_{\mu_{i}}(\mathbf{x}_{t})\geq 0$ for $t=a,a+1,\dots,b$, commonly referred to as “safety” constraints. We employ SafeDiffuser Xiao et al. (2023) to generate trajectories that comply with these safety constraints. SafeDiffuser integrates control barrier functions (CBFs) Nguyen and Sreenath (2016) to enforce finite-time diffusion invariance directly within the sampling process.
4 Case Study
Figure 3: Planned Trajectory (left) and Actual Execution Trajectory (right) in Case Study. The numbers next to start point and regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ indicate the completion times for each reachability progress assigned by progress allocation module.
Type Details
$\mathbb{P}^{\mathcal{R}}$ $\mathcal{R}(\lambda_{1},\lambda_{1},\mu_{1})$, $\mathcal{R}(\lambda_{1}+\lambda_{2},\lambda_{1}+\lambda_{2},\mu_{2})$,
$\mathcal{R}(\lambda_{1}+\lambda_{2}+\lambda_{3},\lambda_{1}+\lambda_{2}+\lambda_{3},\mu_{3})$,
$\mathcal{R}(0,0,\neg\mu_{4})$, $\mathcal{R}(0,0,\neg\mu_{5})$
$\mathbb{P}^{\mathcal{I}}$ $\mathcal{I}(1,110,\neg\mu_{4})$, $\mathcal{I}(1,110,\neg\mu_{5})$
$\mathbb{T}$ $\lambda_{1}\in[0,35]$, $\lambda_{2}\in[35,45]$, $\lambda_{3}\in[10,30]$
Table 1: STL Task Decomposition Results
To illustrate the workflow of our algorithm, we consider a sequential visit and region avoidance task in the Maze2D (Large) environment Fu et al. (2020).
In this scenario, the agent starts at the yellow point shown in Figure 3 and aims to complete the following STL task:
$\mathrm{F}_{[0,35]}(\mu_{1}\wedge(\mathrm{F}_{[35,45]}(\mu_{2}\wedge\mathrm{F}_{[10,30]}\mu_{3})))\wedge\mathrm{G}_{[0,110]}(\neg\mu_{4}\wedge\neg\mu_{5}),$
where the predicate $\mu_{i}$ represents “reach region $\mu_{i}$,” and $\neg\mu_{i}$ denotes “avoid region $\mu_{i}$”. Intuitively, this STL task requires the agent to sequentially visit circular regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ within specific time intervals while avoiding regions $\mu_{4}$ and $\mu_{5}$ throughout the entire episode. The environment layout and target regions are depicted in Figure 3. Our method leverages only an STL task-agnostic trajectory dataset to train the diffusion model, without prior knowledge of the map or system dynamics.
The decomposed progresses and time constraints are summarized in Table 1. The planning algorithm then assigns completion times to each reachability progress, as indicated by the numbers next to start point and regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ in Figure 3, resulting in a sequence of waypoints with corresponding times. The diffusion model is then used to sequentially generate trajectories between adjacent waypoints while ensuring all invariance progresses are satisfied. The final generated trajectory is shown in the left subfigure of Figure 3.
Finally, to evaluate the feasibility of the planned trajectory, we employ a simple model-free PD controller to track it. The resulting execution trajectory is shown in the right subfigure of Figure 3. Using the open-source library stlpy Kurtz and Lin (2022), we calculate the robustness values for both the planned and actual execution trajectories as 0.180 and 0.115, respectively. Since both values are positive, the trajectories satisfy the STL task requirements.
5 Experiments
Type STL Templates
1 $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge\mathrm{G}(\neg\mu_{2})$
2 $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge\mathrm{F}_{\mathrm{I}_{2}}\mu_{2}$
3 $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$
4 $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}\wedge\mathrm{F}_{\mathrm{I}_{4}}(\mu_{4})))))$
5 $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}))))\wedge\mathrm{G}(\neg\mu_{4})$
6 $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1})\wedge\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2})\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3})\wedge\mathrm{G}(\neg\mu_{4})$
7 $\mathrm{F}_{\mathrm{I}_{1}}(\mathrm{G}_{\mathrm{I}_{2}}(\mu_{1}))\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{2})\wedge\mathrm{G}(\neg\mu_{3})$
8 $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge\mathrm{F}_{\mathrm{I}_{2}}(\mathrm{G}_{\mathrm{I}_{3}}(\mu_{2})))$
9 $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2})\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3})\wedge\mathrm{G}_{\mathrm{I}_{4}}(\mu_{4}))$
Table 2: STL Task Templates for Experiments
To evaluate the performance of our algorithm, we further conduct experiments in the Maze2D environment. Specifically, to validate the zero-shot generalization capability of our algorithm for STL tasks, we test it on a set of testing cases containing randomly generated STL tasks in three different Maze2D environments: U-Maze, Medium, Large. The experimental setup follows the same framework described in the Case Study. The agent starts from a randomly generated position and must complete the randomly generated STL tasks by reaching the target region within the specified time interval. All experiments were conducted on a PC running Ubuntu 22.04, equipped with an Intel i7-13700K CPU and an Nvidia 4090 GPU.
STL Tasks Generation
In order to generate random STL tasks, we design nine STL task templates as shown in Table 2. For each template, we randomly generated time intervals and the positions and sizes of circular regions corresponding to atomic predicates in the template, resulting in randomized STL tasks. Specifically, for each pair of Maze2D environment (U-Maze, Medium, Large) and each task template listed in Table 2, we generate 150 feasible random STL formulae.
Baseline Algorithm
We compare our algorithm with the method proposed in Zhong et al. (2023), which adopts classifier-based guidance and directly leverages the gradient of the trajectory’s robustness value to guide the sampling process of the diffusion model, thereby optimizing the robustness of the generated trajectory. The gradient of robustness is calculated by the STLCG method proposed in Leung et al. (2023). In the following text, we refer to this algorithm as the Robustness Guided Diffuser (RGD).
Experiment Settings
The diffusion models used in both RGD and our algorithm are trained following the procedure in Janner et al. (2022) using the D4RL dataset Fu et al. (2020). A simple multilayer perceptron (MLP) with four fully connected layers is used as the TimePredict model in our algorithm and it is also trained using the D4RL dataset. In our experiments, we employ diffusion model to generate only the state sequence of the trajectory and use a simple PD controller to follow the state sequence during running to get the actual execution trajectory.
Evaluation Metrics
For each environment and each STL task template, we test RGD and our algorithm on all randomly generated test cases and record the average of the following metrics across all cases:
  • Execution Success Rate (SR): The proportion of cases where the actual execution trajectory achieve non-negative robustness values.
  • Total Planning Time (T0): The average total running time (in seconds) to plan a trajectory per case.
In addition, we also record the average Trajectory Generation Time (T1), which is the average time spent by the Trajectory Generation module of our algorithm per case. By recording this metric, we analyze the proportion of runtime contributed by each module in our algorithm.
Env Type Success Rate(%$\uparrow$ Total Planning Time(s)$\downarrow$ T1(s)
RGD ours RGD ours
U 1 80.00 97.33 13.43$\pm$1.51 0.86$\pm$0.13 0.86
2 36.67 92.00 16.65$\pm$2.06 0.64$\pm$0.17 0.64
3 32.00 91.33 19.68$\pm$11.76 1.31$\pm$0.15 1.21
4 - 90.00 - 1.64$\pm$0.20 1.38
5 - 84.67 - 2.59$\pm$0.45 2.46
6 - 86.67 - 2.35$\pm$0.42 2.35
7 - 89.33 - 1.86$\pm$0.33 1.86
8 - 97.33 - 0.98$\pm$0.15 0.81
9 - 88.67 - 1.53$\pm$0.27 1.52
M 1 70.00 94.67 53.90$\pm$5.78 3.74$\pm$0.34 3.74
2 34.67 89.33 70.69$\pm$8.27 2.72$\pm$0.67 2.72
3 35.33 83.33 129.87$\pm$27.25 5.53$\pm$0.33 5.43
4 - 83.33 - 7.09$\pm$0.54 6.80
5 - 82.00 - 11.36$\pm$1.32 11.22
6 - 84.67 - 11.76$\pm$1.36 11.76
7 - 90.00 - 8.01$\pm$1.21 8.00
8 - 91.33 - 3.87$\pm$0.38 3.71
9 - 82.67 - 6.53$\pm$0.74 6.52
L 1 34.67 92.00 55.48$\pm$5.31 3.62$\pm$0.33 3.62
2 16.67 81.33 68.70$\pm$7.65 2.88$\pm$0.63 2.87
3 26.67 79.33 136.35$\pm$38.12 5.59$\pm$0.34 5.49
4 - 69.33 - 7.46$\pm$0.52 7.13
5 - 79.33 - 12.62$\pm$0.75 12.45
6 - 73.33 - 12.37$\pm$1.35 12.37
7 - 84.00 - 8.18$\pm$0.68 8.18
8 - 85.33 - 3.93$\pm$0.33 3.75
9 - 76.67 - 6.93$\pm$0.60 6.92
Table 3: Result of Experiment in Maze2D Environment. U:U-Maze; M:Medium; L:Large; RGD: Robustness Guided Diffuser; T1:Trajectory Generation Time.
Results and Analysis
The experimental results are shown in Table 3. During testing, we found that the Robustness Guided Diffuser works only for simple STL tasks (Types 1 to 3). For more complex tasks, it faces two key issues: (1) Optimization Challenges: It is often infeasible to achieve a non-negative robustness value within a limited number of denoising steps; and (2) High Computational Cost: Complex tasks require longer trajectory generations and more complex robustness value computations, significantly increasing the computational overhead. As a result, the Robustness Guided Diffuser is suitable only for simple, short-horizon STL tasks.
In contrast, our method decomposes complex STL tasks into several short-horizon trajectory generation tasks under simpler constraints, significantly improving both planning success rate and efficiency. In relatively simple environments such as U-Maze and Medium, our algorithm achieves an actual execution success rate of over 80% across all types of STL tasks. Even in the more complex environment, Large, our algorithm maintains a success rate of at least 69% for all task templates. Additionally, the total planning time of our method is significantly lower than that of the Robustness Guided Diffuser (more than 10× faster).
By comparing the Trajectory Generation Time (T1) and Total Planning Time (T0), we identify trajectory generation as the primary efficiency bottleneck of our algorithm. For tasks with more predicates (Type 4, 9) or those involving “avoid” predicates (Type 3, 5, 6 and 7), runtime increases accordingly. This is mainly because tasks involving more predicates require longer waypoint sequences to satisfy, necessitating multiple calls to the diffusion model for trajectory generation. Additionally, for tasks involving “avoid” predicates, the use of SafeDiffuser requires solving more complex quadratic programming (QP) problems to ensure the trajectories satisfy the invariance progresses, thereby increasing the runtime. Recent advancements in accelerating the sampling process of Diffusion Models Song et al. (2020); Xiao et al. (2021); Zhou et al. (2025) could further enhance the efficiency of our approach, which will be studied in our future work.
Further Experiments
In the above experiments, since the system model in the simulation environment is unknown, we cannot rely on model-based approaches to precisely determine the feasibility of the random generated tasks. Instead, the feasibility of the STL formulas is assessed using our progress allocation module without considering the trajectory generation module. This approach may, to some extent, lead to an optimistic estimation of the success rate of our algorithm. To further evaluate the planning success rate, we utilize a custom-built environment where we have full access to both the environment information and system dynamics. In this setup, we use an optimization-based algorithm as a sound and complete solution Gilpin et al. (2020) to accurately determine the feasibility of the STL formulas. Experimental results demonstrate that the progress allocation module in our algorithm achieves a success rate of over 80% across various task scenarios, underscoring its strong completeness and robustness. Details of these additional experiments are provided in the Appendix C.
6 Conclusion
This paper introduced a hierarchical framework for trajectory planning under Signal Temporal Logic (STL) constraints. By integrating task decomposition with a diffusion model pre-trained on task-agnostic data, the proposed approach achieved high success rates, scalability, and zero-shot generalization to various STL tasks, while significantly reducing computational costs comparing to baseline algorithm. In future work, we aim to further optimize the progress allocation module by introducing uncertainty-aware time prediction and iterative optimization mechanism. These enhancements are expected to reduce the conservativeness of the algorithm and improve the solution quality. Additionally, we plan to integrate receding horizon control and accelerated diffusion sampling to further enhance the execution success rate and improve the algorithm’s efficiency.
Appendix A Related Works
A.1 STL Decomposition
To address the high complexity of STL control synthesis problems, several decomposition-based methods have been proposed Leahy et al. [2023]; Yu et al. [2023]; Kapoor et al. [2024]. In Leahy et al. [2023], the authors proposed a formula transformation-based method for multi-agent STL planning. This approach jointly decomposes an STL specification and team of agents. In Yu et al. [2023], the authors decompose STL tasks into several subtasks with non-overlapping time intervals using time interval decomposition and sequentially apply the shrinking horizon Model Predictive Control (MPC) algorithm to each short-time-interval subtask. However, this method is limited to handling STL fragments that do not include nested temporal logic operators.
The work most similar to our STL decomposition framework is Kapoor et al. [2024]. This work first decomposes STL tasks into several spatio-temporal subtasks with time variable constraints. Then, through time variable simplification, partial ordering, and slicing, the subtasks are segmented into several time intervals. Finally, a planning algorithm is subsequently used to sequentially solve the atomic tasks within each time interval. Our algorithm similarly begins by decomposing the STL task into several spatio-temporal progresses and time variable constraints. However, we adopt a search-based approach to determine the completion times and corresponding states to achieve each progress. During the search process, our method dynamically maintains the time variable constraints on-the-fly according to the completion of progresses. By incorporating the search mechanism, our algorithm achieves greater completeness compared to the incremental planning approach used in Kapoor et al. [2024]. Additionally, the dynamic maintenance of time variable constraints enables a more natural handling of relationships between subtasks and extends the applicability of our approach to more complex STL task fragments that allows “until” operator.
A.2 Planning with Diffusion Model
Recent advancements in diffusion-based planning methods highlight their remarkable flexibility, as they rely exclusively on offline trajectory datasets and do not require direct interaction with or access to the environment. By leveraging guided sampling, these methods can address a wide range of objectives without the need for retraining. This approach has been widely applied to long-horizon task planning and decision-making, facilitating the generation of states or actions for control purposes Janner et al. [2022]; Ajay et al. [2022]; Chi et al. [2023].
In the domain of diffusion-based planning for temporal logic tasks, several significant studies have been conducted Zhong et al. [2023]; Meng and Fan [2024]; Feng et al. [2024a, b]. For instance, Feng et al. [2024a] proposed a classifier-based guidance approach to direct the sampling process of the diffusion model, enabling the generation of trajectories that fulfill finite Linear Temporal Logic ($\text{LTL}_{f}$) tasks. Similarly, Feng et al. [2024b] introduced a data-driven hierarchical framework that decomposes co-safe LTL tasks into sub-tasks using hierarchical reinforcement learning. This framework integrates a diffusion model with a determinant-based sampling technique to efficiently produce diverse low-level trajectories, enhancing both planning success rates and task generalization capabilities.
For diffusion-based Signal Temporal Logic (STL) planning, Zhong et al. [2023] employed a classifier-based guidance method that leverages the gradient of robustness values to guide the sampling process of a diffusion model. This method enabled the generation of vehicle trajectories compliant with STL-specified traffic rules. Expanding on this work, Meng and Fan [2024] introduced a data augmentation process to further improve trajectory diversity and increase the satisfaction rate of specified rules. However, these approaches remain constrained to relatively simple STL tasks due to the inherent complexity of optimizing robustness values, as well as the trade-off between maximizing reward objectives and preserving the feasibility of generated trajectories Li et al. [2024].
We compare our algorithm with the method proposed in Zhong et al. [2023] in the experiments. The results demonstrate that our algorithm can handle complex and long-horizon STL tasks that the method in Zhong et al. [2023] cannot address, while also exhibiting significant advantages in runtime efficiency.
Appendix B Diffusion Models for Trajectory Planning
The core concept of diffusion-model-based trajectory planningJanner et al. [2022]; Ajay et al. [2022] is to employ the diffusion model to learn the distribution of pre-collected trajectories $q\left(\boldsymbol{\tau}^{0}\right)$ in the current environment under the system dynamics, thereby transforming planning or control synthesis problems into conditional trajectory generation.
Diffusion models consist of two processes: the diffusion process and the denoising process.
The diffusion process gradually adds Gaussian noise to a trajectory $\boldsymbol{\tau}^{0}$, transforming it into noise. At each timestep $i$, the noisy trajectory is given by:
$q(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i-1})=\mathcal{N}(\boldsymbol{\tau}^{i};\sqrt{1-\beta_{i}}\boldsymbol{\tau}^{i-1},\beta_{i}\mathbf{I}),$ (B.1)
where $\beta_{i}$ controls the noise scale, and $N$ is the total number of diffusion steps. The noisy trajectory $\boldsymbol{\tau}^{i}$ at any step can be directly computed as:
$\boldsymbol{\tau}^{i}=\sqrt{\bar{\alpha}_{i}}\boldsymbol{\tau}^{0}+\sqrt{1-\bar{\alpha}_{i}}\boldsymbol{\varepsilon},\quad\boldsymbol{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (B.2)
with $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$.
The denoising process reverses the diffusion by iteratively recovering the trajectory from Gaussian noise. The reverse distribution is approximated as:
$p_{\theta}(\boldsymbol{\tau}^{i-1}\mid\boldsymbol{\tau}^{i})=\mathcal{N}(\boldsymbol{\tau}^{i-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i),\boldsymbol{\Sigma}^{i}),$ (B.3)
where $\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)$ is parameterized by the model, and $\boldsymbol{\Sigma}^{i}$ is typically fixed. Instead of learning $\boldsymbol{\mu}_{\theta}$ directly, the model typically predicts the noise $\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)$ and gets $\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)$ according to the relationshipHo et al. [2020]:
$\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)=\frac{1}{\sqrt{\alpha_{i}}}\left(\boldsymbol{\tau}^{i}-\frac{1-\alpha_{i}}{\sqrt{1-\bar{\alpha}_{i}}}\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)\right).$ (B.4)
The model is trained by minimizing a simplified loss function:
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{i,\boldsymbol{\varepsilon},\boldsymbol{\tau}^{0}}\left[\left\|\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)\right\|_{2}^{2}\right].$ (B.5)
In this paper, superscripts denote the diffusion time step, while subscripts indicate the trajectory time step. For example, $\boldsymbol{\tau}^{i}_{t}$ represents the state at trajectory time step $t$ during diffusion time step $i$. For noise-free trajectories, $\boldsymbol{\tau}^{0}$, we omit the superscript when there is no ambiguity.
B.1 Safe Planning with Diffusion Model
Previous studies Xiao et al. [2023]; Botteghi et al. [2023]; Mizuta and Leung ; Zheng et al. [2024]; Christopher et al. [2024] have demonstrated the effectiveness of diffusion models in generating safety constraint-compliant trajectories. In this work, we employ SafeDiffuser Xiao et al. [2023], which integrates control barrier functions (CBFs) Nguyen and Sreenath [2016] to enforce finite-time diffusion invariance directly within the sampling process.
SafeDiffuser models the denoising process as a dynamic system:
$\dot{\boldsymbol{\tau}}^{i}=\lim_{\Delta\tau\to 0}\frac{\boldsymbol{\tau}^{i}-\boldsymbol{\tau}^{i+1}}{\Delta\boldsymbol{\tau}}=\boldsymbol{u}^{i},$ (B.6)
where $\Delta\boldsymbol{\tau}$ represents the diffusion time step, and $\boldsymbol{u}^{i}$ is a control variable with the same dimension as $\boldsymbol{\tau}^{i}$.
For each constraint $b(\mathbf{x}_{t})=h_{\mu_{i}}(\mathbf{x}_{t})\geq 0$, we define the associated CBF constraints as:
$\frac{\mathrm{d}b\left(\mathbf{x}_{t}^{i}\right)}{\mathrm{d}\mathbf{x}_{t}^{i}}\boldsymbol{u}_{t}^{i}+\boldsymbol{\alpha}(b\left(\mathbf{x}_{t}^{i}\right))\geq 0,\quad\forall i\in\{0,\dots,N-1\},$ (B.7)
where $\boldsymbol{\alpha}$ is an extended class-$\mathcal{K}$ function, and $N$ is the total number of diffusion time steps.
To ensure compliance with these constraints while minimally altering the denoising process, the optimal control $\boldsymbol{u}^{i*}$ is computed at each diffusion time step $i$ by solving the following quadratic programming (QP) problem:
$\text{minimize}\quad\left\|\boldsymbol{u}^{i}-\frac{\boldsymbol{\tau}^{i}-\boldsymbol{\tau}^{i+1}}{\Delta\boldsymbol{\tau}}\right\|^{2},$ (B.8)
subject $\text{subject to}\quad\frac{\mathrm{d}b\left(\mathbf{x}_{t}^{i}\right)}{\mathrm{d}\mathbf{x}_{t}^{i}}\boldsymbol{u}_{t}^{i}+\boldsymbol{\alpha}(b\left(\mathbf{x}_{t}^{i}\right))\geq 0.$
The updated state $\boldsymbol{\tau}^{i*}$ is then computed by substituting $\boldsymbol{u}^{i}=\boldsymbol{u}^{i*}$ into (B.6), and $\boldsymbol{\tau}^{i}$ is replaced with $\boldsymbol{\tau}^{i*}$. The complete denoising procedure is summarized in Algorithm 3.
Algorithm 3 Safe Planning with Diffusion Model
0:  Trajectory length $T$, start state $\mathbf{x}_{s}$, goal state $\mathbf{x}_{g}$, invariance constraints $\{b(\mathbf{x}_{t})\geq 0\}$
0:  A trajectory $\boldsymbol{\tau}$ of length $T$ that satisfies all constraints
1:  Initialize $\boldsymbol{\tau}^{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ with length $T$
2:  Replace $\mathbf{x}_{0},\mathbf{x}_{T}$ in $\boldsymbol{\tau}^{N}$ with $\mathbf{x}_{s},\mathbf{x}_{g}$
3:  for $i=N-1,...,0$ do
4:     $\boldsymbol{\mu}^{i+1}=\frac{1}{\sqrt{\alpha_{i+1}}}(\boldsymbol{\tau}^{i+1}-\frac{1-\alpha_{i+1}}{\sqrt{1-\bar{\alpha}_{i+1}}}\boldsymbol{\varepsilon_{\theta}}(\boldsymbol{\tau}^{i+1},i+1))$
5:     $\boldsymbol{\tau}^{i}\sim\mathcal{N}(\boldsymbol{\mu}^{i+1},\boldsymbol{\Sigma}^{i+1})$
6:     Solve the QP (B.8) and get $\boldsymbol{u}^{i*}$
7:     Calculate $\boldsymbol{\tau}^{i*}$ by setting $\boldsymbol{u}^{i}=\boldsymbol{u}^{i*}$ in (B.6)
8:     $\boldsymbol{\tau}^{i}\leftarrow\boldsymbol{\tau}^{i*}$
9:     Replace $\mathbf{x}_{0},\mathbf{x}_{T}$ in $\boldsymbol{\tau}^{i}$ with $\mathbf{x}_{s},\mathbf{x}_{g}$
10:  end for
11:  return  $\boldsymbol{\tau}^{0}$
B.2 Time-aware Trajectory Planning
As noted in Janner et al. [2022], the planning horizon of a diffusion model is not fixed by its architecture but dynamically adapts based on the input noise size. This allows for variable-length plans during the denoising process at test time. In this work, we leverage this adaptability by directly adjusting the input noise size to generate trajectories of the desired length.
During experiments, we observe that models trained with longer planning horizons generalize better to different trajectory lengths during testing. In contrast, models trained with shorter planning horizons exhibit weaker generalization. To address this, we recommend either:
  • Using variable horizons during training instead of a fixed horizon, or
  • Employing multiple models trained with different horizons to generate trajectories of diverse lengths.
Appendix C Comparative Experiment with Optimization-based Method
To further evaluate the success rate of the progress allocation module in our algorithm, we compare it against a widely-used optimization-based algorithm Gilpin et al. [2020] in a custom-built simulation environment. The baseline algorithm is employed as a sound and complete solution to accurately assess the feasibility of randomly generated test cases.
The experiment is conducted within a bounded $10\times 10$ square 2D plane containing a circular obstacle. The underlying system dynamics are modeled using a double integrator. The agent starts from a randomly generated position and must complete the randomly generated STL tasks by reaching the target region within the specified time interval.
In this experiment, the baseline algorithm is implemented using the open-source library stlpy Kurtz and Lin [2022] and has full knowledge of the environmental information and system dynamics, while our algorithm only has access to the trajectory dataset.
To generate the trajectory dataset, we randomly sample start and end points in the environment and use the baseline algorithm to solve reach-avoid tasks. This process produces 200,000 collision-free trajectories that satisfy the system dynamics, which are then used to train the diffusion model.
Following the training procedure from Janner et al. [2022], we train two diffusion models with different planning horizons, as described in Section B.2. The first model is trained on trajectory segments of length 16 for short trajectories, while the second model uses segments of length 32 to improve generalization to longer trajectories.
We generate 200 feasible STL tasks for each template, as described in Section 5. The deterministic baseline algorithm is used to ensure the feasibility of these tasks. However, for templates 4 and 5, which involve multi-layer nesting of temporal operators, the baseline algorithm fails to find solutions within an acceptable time. In these cases, we still employ our algorithm’s progress allocation module to verify feasibility.
In addition to the Execution Success Rate (SR) and Total Planning Time (T0) metrics described in Section 5, we introduce an additional evaluation metric:
  • Progress Allocation Success Rate (SR0): The proportion of cases where the progress allocation module successfully identifies a sequence of waypoints. This metric specifically measures the reliability of the progress allocation module in our algorithm.
Type SR0(%) SR(%) Total Planning Time(s)$\downarrow$
ours baseline
1 96.0 93.5 0.99$\pm$0.08 3.82$\pm$1.44
2 98.0 96.5 0.81$\pm$0.03 6.30$\pm$1.36
3 96.0 89.0 1.26$\pm$0.05 31.60$\pm$10.46
4 - 78.5 2.10$\pm$0.26 Timeout
5 - 83.0 2.57$\pm$0.10 Timeout
6 97.5 69.5 2.87$\pm$0.12 24.23$\pm$6.39
7 80.0 73.5 1.80$\pm$0.08 7.71$\pm$3.50
8 89.5 89.0 0.82$\pm$0.03 106.58$\pm$82.19
9 81.0 72.0 1.61$\pm$0.06 151.19$\pm$78.82
Table C.1: Result of Experiment in Custom-built Environment. SR0: Progress Allocation Success Rate; SR: Execution Success Rate;
The experimental results are summarized in Table C.1. Our algorithm achieves consistently high success rates across test cases generated from diverse task templates. Notably, the Execution Success Rate (SR) exceeds 70% in all scenarios, demonstrating the algorithm’s strong generalization capability for STL tasks.
For all templates except 4 and 5, the Progress Allocation Success Rate (SR0) exceeds 80%, indicating that the progress allocation module is generally reliable, albeit slightly conservative.
Finally, by comparing the Total Planning Time, our algorithm significantly outperforms the optimization-based baseline algorithm, highlighting the efficiency of the task decomposition and planning framework employed in our approach. Notably, for templates 4 and 5, which involve multi-layered nested STL tasks, the baseline algorithm fails to find feasible solutions within a reasonable time. In contrast, our algorithm demonstrates both high success rates and high efficiency, even in these complex scenarios.
Appendix D Implementation Details of Experiments
D.1 Trajectory Generation
In this work, we employ diffusion model to generate only the state sequence of the trajectory to avoid the practical issues discussed in Ajay et al. [2022]. During testing, we use a simple PD controller to follow the state sequence to get the action sequence on-the-fly and generate the final actual execution trajectory, as it performs better than employing an inverse dynamics model used in Ajay et al. [2022] to generate actions.
D.2 Calculation of the Robustness Value
Besides the Boolean semantics, STL also incorporates a concept of robustness, which refers to its quantitative semantics used to measure the degree to which a signal satisfies or violates a formula. Positive robustness values signify satisfaction, while negative values indicate violation. The quantitative semantics of a STL formula $\varphi$ with respect to a signal $\mathbf{s}_{t}$ starting at time $t$ are defined as follows:
$\rho\left(\mathbf{s}_{t},\top\right)$ $=\rho_{\max},\text{ where }\rho_{\max}>0,$ (D.1)
$\rho\left(\mathbf{s}_{t},\mu\right)$ $=h_{\mu}(\mathbf{x}_{t}),$
$\rho\left(\mathbf{s}_{t},\neg\mu\right)$ $=-h_{\mu}(\mathbf{x}_{t}),$
$\rho\left(\mathbf{s}_{t},\varphi_{1}\wedge\varphi_{2}\right)$ $=\min\left(\rho\left(\mathbf{s}_{t},\varphi_{1}\right),\rho\left(\mathbf{s}_{t},\varphi_{2}\right)\right),$
$\rho\left(\mathbf{s}_{t},\varphi_{1}\vee\varphi_{2}\right)$ $=\max\left(\rho\left(\mathbf{s}_{t},\varphi_{1}\right),\rho\left(\mathbf{s}_{t},\varphi_{2}\right)\right),$
$\rho\left(\mathbf{s}_{t},\mathrm{F}_{[a,b]}\varphi\right)$ $=\max_{t^{\prime}\in[t+a,t+b]}\rho\left(\mathbf{s}_{t^{\prime}},\varphi\right),$
$\rho\left(\mathbf{s}_{t},\mathrm{G}_{[a,b]}\varphi\right)$ $=\min_{t^{\prime}\in[t+a,t+b]}\rho\left(\mathbf{s}_{t^{\prime}},\varphi\right),$
$\rho\left(\mathbf{s}_{t},\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2}\right)$ $=\max_{t^{\prime}\in[t+a,t+b]}\min\{\rho\left(\mathbf{s}_{t^{\prime}},\varphi_{2}\right),$
$\quad\quad\quad\quad\quad\quad\min_{\tau\in\left[t,t^{\prime}\right]}\rho\left(\mathbf{s}_{\tau},\varphi_{1}\right)\}.$
In the experiments, we utilized the commonly adopted open-source library stlpy Kurtz and Lin [2022] to compute the robustness values of both the planned trajectories and the actual execution trajectories. In the Maze2D environment, transitions between states require relatively long trajectories, which result in extended time intervals for the corresponding STL tasks. This creates computational overhead for calculating robustness values using stlpy, exceeding the capacity of the experimental hardware.
To address this issue, we introduced an additional sampling factor, $\eta$, which represents the mapping relationship between the time length in the STL task and the actual system trajectory. Specifically, we assume that each time step in the STL task corresponds to a trajectory length of $\eta$ in the system. When computing robustness values, we sample one state from the system trajectory for every $\eta$ states, resulting in a sampled trajectory. The robustness value is then calculated based on this sampled trajectory. While this operation may introduce minor deviations in the robustness value calculations, it does not affect the validity of the experimental conclusions.
D.3 Experimental Parameter Settings
Some of the parameters involved in the experiments are listed below, and their specific values are shown in Table D.1:
  • Maximum Number of Attempts ($N_{\mathrm{max}}$): The maximum number of attempts for new state sampling in Algorithm 2.
  • Horizon ($H$): The planning horizon used during the training of the diffusion model.
  • Total Denoise Steps ($N$): The total number of steps in the denoising process.
  • Scaling Factor ($\gamma$): Applied to the predicted mean trajectory length, used to control the conservativeness of the Progress Allocation Module, as described in Section 3.3.
  • Sampling Factor ($\eta$): Used when computing the robustness value of trajectories, as described in Section D.2.
Env $N_{\mathrm{max}}$ $H$ $N$ $\gamma$ $\eta$
Maze2D-Umaze 1 128 64 0.8 8
Maze2D-Medium 1 256 256 0.9 8
Maze2D-Large 1 384 256 1.1 8
Custom-Built Env 1 16&32 64 1 1
Table D.1: Parameters Used in the Experiments
Appendix E More Cases
We visualized some of the experimental results. The actual execution trajectories for some successful cases (where the STL tasks were satisfied by the execution trajectories) are shown in Figure E.1 to Figure E.3. For some failed cases (where the STL tasks were not satisfied), the trajectories planned by our algorithm are shown in Figure E.4.
Notably, by analyzing the failed cases, we identified that the primary reason for execution failure is that the trajectories generated by the diffusion model significantly violated system dynamics, such as colliding with obstacles in the environment or having excessively large distances between consecutive states. To further enhance the actual execution success rate, our method can be integrated with some receding horizon control methods Zhou et al. [2024a] or online replanning strategies Zhou et al. [2024b]. This extension will be explored in our future work.
Figure E.1: Successful Execution Trajectories in Maze2D-large Environment under STL Template 3: $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$. The task requires the agent to eventually reach region $\mu_{1}$ within a given time interval, but before reaching $\mu_{1}$, it must first visit region $\mu_{2}$.
Figure E.2: Successful Execution Trajectories in Maze2D-medium Environment under STL Template 4: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}\wedge\mathrm{F}_{\mathrm{I}_{4}}(\mu_{4})))))$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3},\mu_{4}$ within the given time intervals.
Figure E.3: Successful Execution Trajectories in Maze2D-umaze Environment under STL Template 5: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}))))\wedge\mathrm{G}(\neg\mu_{4})$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3}$ within the given time intervals and always avoid the region $\mu_{4}$.
Figure E.4: Planned Trajectories in Some Failure Cases.